# Privy: Envisioning and Mitigating Privacy Risks for Consumer-facing AI Product Concepts

### v1

@@ 5.3. Data Analysis
We applied a mixed-methods approach to analyzing our data. Quantitative analysis. We calculated descriptive statistics (means, standard deviations) to examine how Privy-LLM and Privy-Template influenced practitioners’ ability to identify and mitigate privacy risks (RQ1). We fit ordinal logistic regressions to test how Privy version affected (i) privacy experts’ quality ratings of reports and (ii) practitioners’ self-reported usefulness of Privy in risk envisioning and mitigation (RQ2). Privy version was the main predictor (Privy-Template as reference). For expert evaluations, we included random intercepts for expert ID and, where applicable, participant ID to account for repeated measures. [^4] Specifically, for the “Per Risk” quality measure in Table 3, we included random intercepts for both expert ID and participant ID.
@@ 6. Evaluation User Study Findings
6.1. RQ1: How and to what extent do Privy-LLM and Privy-Template help practitioners identify and mitigate privacy risks with AI product concepts?
We first analyze how Privy, which in both versions helps foreground privacy considerations in designing AI products, shapes practitioners’ approaches to risk identification (Section 6.1.1) and risk mitigation (Section 6.1.2). Since the findings reported in this section are shared across both Privy-LLM and Privy-Template, we refer to them collectively as “Privy” for simplicity.
@@ 6.1.1. Privy helps practitioners envision privacy risks that are relevant and severe
Our findings suggest that Privy helped participants identify privacy risks that were both relevant and severe to their AI product concepts (DG1, DG2), from both the practitioners’ own and the privacy experts’ perspectives. Specifically, participants rated 55 (76%) of the 72 privacy risks they themselves identified as being of high relevancy and 47 (65%) of high severity. Similarly, on a 6-point Likert scale (6 as strongly agree), privacy experts, on average, rated these risks as 4.58 (SD=1.32) on relevancy and 4.97 (SD=1.07) on severity (Table 4).
@@
Privy (Overall) Privy-LLM Privy-Template
@@
  • *Note. 1: strongly disagree, 2: disagree, 3: slightly disagree, 4: slightly agree, 5: agree, 6: strongly agree
@@ 6.1.2. Privy helps practitioners brainstorm risk mitigation plans that are effective and appropriate
Taken together, these results show that Privy enabled non-experts to articulate risk mitigation plans that could initiate broader privacy conversations, though additional support is needed to generate mitigation strategies that are immediately actionable and effective — a challenge we return to in Section 6.3.
This analysis also points to a notable gap in expert-assessed quality between the plans produced with Privy-LLM versus Privy-Template (Table 5). We explore the effects of the use of LLMs on the quality of privacy impact assessments in Section 6.2.
@@
Privy (Overall) Privy-LLM Privy-Template
@@
  • *Note. 1: strongly disagree, 2: disagree, 3: slightly disagree, 4: slightly agree, 5: agree, 6: strongly agree
@@
We next examined differences between Privy-LLM and Privy-Template. First, we compare how privacy experts rated the quality of the privacy impact assessments produced by both Privy-LLM versus Privy-Template (Section 6.2.1). Then, we explore how practitioners perceived the usefulness of Privy-LLM versus Privy-Template in helping them complete their privacy impact assessments (Section 6.2.2). In both cases, we complement our quantitative analysis with qualitative insights into participants’ usage patterns and opinions towards Privy’s LLM-powered features (Section 6.2.3).
6.2.1. The use of LLMs improves the quality of privacy impact assessments produced by Privy
We evaluated the quality of privacy impact assessments produced with Privy-LLM and Privy-Template by analyzing the privacy risks identified and mitigation plans proposed by 24 participants (Privy-LLM: $N=12$; Privy-Template: $N=12$). Thirteen privacy experts independently rated these outputs on multiple quality measures (Fig. 6). Our ordinal logistic regression analyses revealed that participants using Privy-LLM produced risk statements rated significantly higher in relevancy ($\beta=1.95,p<.01$), severity ($\beta=2.13,p<.001$), correctness ($\beta=2.10,p<.001$), and clarity ($\beta=2.39,p<.001$), compared to those generated with Privy-Template. A similar pattern emerged for risk mitigation plans: experts judged the Privy-LLM group’s mitigation plans to be significantly more effective — “addressing identified risks” ($\beta=1.76,p<.001$) and “useful conversation starter” ($\beta=2.51,p<.001$) — and appropriate — “product specificity” ($\beta=2.04,p<.001$) and “practicality” ($\beta=1.65,p<.01$). Overall, Privy-LLM significantly outperformed Privy-Template across all quality measures, including overall risk identification ($\beta=2.45,p<.001$) and mitigation ($\beta=1.95,p<.001$) quality, as well as the quality of the privacy impact assessment as a whole ($\beta=2.45,p<.001$) (Fig. 7).
@@ 6.2.2. The use of LLMs has minimal, if any, impact on the perceived usefulness and usability of Privy
Beyond assessing output quality, we compared Privy-LLM against Privy-Template on practitioners’ perceived usefulness and overall usability ratings across our five design goals (DG1-DG5). Both versions scored highly across these measures, indicating that each effectively helped users in achieving the intended goals (Fig. 8). System Usability Scale (SUS) scores were similarly strong for Privy-LLM (M=76.04, SD=15.65) and Privy-Template (M=75.63, SD=18.19), reflecting a “good” overall level of usability (Bangor et al., 2009).
Figure 8. Participants perceived that both Privy-LLM and Privy-Template were useful in achieving the design goals (DG1-DG5).
An ordinal logistic regression showed no significant differences between versions on perceived usefulness, and a linear regression on SUS scores likewise showed no significant difference. Taken together with Section 6.2.1, these findings suggest that while Privy-LLM supports the production of higher-quality privacy impact assessments, practitioners view both versions of Privy as comparably useful and usable.
@@ 6.2.3. Use patterns and concerns with human-AI collaboration in privacy risk envisioning and mitigation
We next explored how participants used the LLM-powered features in Privy-LLM, and their concerns therein.
@@
Some participants used Privy-LLM as a thought partner, offering validation and reassurance. For example, as P3 put it, it served as a sanity check on their reasoning when working alone. P9 even described an “aha moment” when an LLM suggestion mirrored their own thinking: “I had come up with something without having fully read the AI suggestion that the AI thought like, this is actually something you may want to consider.”
@@
Despite the benefits of LLM support, participants raised two main concerns: overestimating the level of automated assistance provided by Privy-LLM, and the risk of overreliance on its outputs. These concerns point to a core tension: while LLMs enhanced privacy assessment quality (Section 6.2.1), they also challenged practitioners’ sense of agency in privacy work.
Some participants felt that parts of the workflow in Privy-LLM, e.g., risk mitigation, were still driven primarily by their own effort and could be provided with more LLM-assistance. As P7 noted, “I think the brainstorming mitigation tasks, I came up with them… I don’t think I took so much help from Privy.” Others described the extra cognitive effort required to integrate AI suggestions into their privacy workflow: there’s 7–12 different things [Privy-LLM-suggested risks]… that’s a lot to juggle in my mind” (P9).
Participants also worried that LLM-generated content could bias their thinking and diminish cognitive engagement. For example, P7 made a point to generate their own ideas before consulting Privy-LLM’s provocations: “just making sure that doesn’t create bias in my head.” Similarly, P1 emphasized pausing to reflect: “have I included the risks that I felt were the most important ones before I started looking at the suggestions?” P9 further warned that less experienced practitioners may be more prone to overreliance and uncritical acceptance: “you may actually just be like, sure, this looks good, and accept.”
@@
In post-activity interviews, participants reflected on key challenges in envisioning and mitigating privacy risks for AI product concepts. We first synthesize these challenges (Section 6.3.1), drawing on the SPAF framework (Das et al., 2022) to categorize them into barriers of awareness, motivation, and ability. We then examine how Privy — across both Privy-LLM and Privy-Template — helped participants address some of these barriers (Section 6.3.2), and where limitations remain (Section 6.3.3).
@@
AI products are often highly complex, making it difficult to develop a comprehensive view of privacy risks. Practitioners described the absence of a systematic risk scanning mechanism, particularly when products rely on multiple data flows and third-party integrations. As P20 explained, a key challenge was “figuring out all the types of data you are taking from the user… but then you’re using some third-party solution in your product, and they’re using it, and you’re not aware of it.”
@@
Echoing prior work (Lee et al., 2024a), participants stressed that privacy is rarely treated as a core priority. P22 anticipated organizational resistance even if Privy were adopted: “I can already imagine my boss will say… this is not our first priority. We don’t need to worry about that… only if someone brings it up, or if the organization has a concern.” Some participants also described a tendency to externalize responsibility, framing privacy as the domain of specialized teams rather than part of their own remit. This distancing reduced ownership: unless privacy experts explicitly raised concerns, practitioners felt little obligation. As P6 noted, “when these external teams don’t tell us [about privacy issues], then we’re not responsible for it.”
@@
6.3.2. Privy-LLM and Privy-Template address major barriers
We found that Privy-LLM and Privy-Template could help practitioners address these barriers.
@@ 6.3.2. Privy-LLM and Privy-Template address major barriers
Privy’s structured workflow provided critical scaffolding for envisioning privacy risks. Participants appreciated how the tool’s concrete definition of privacy risks and taxonomy anchored their reasoning: “understand the taxonomy here and then apply this taxonomy” (P18). Others highlighted how the workflow supported prioritization, helping them discern underlying issues. As P17 reflected, “comparing [risks] was really fun… I started to realize, oh, maybe surveillance is not the big issue. Insecurity is the underlying cause.”
Privy’s structured workflow enables practitioners to ground risk envisioning from various aspects (Section 6.1.1) and to reduce blind spots (Section 6.2.3), even without LLM-powered features. As P21 noted after using Privy-Template: “there was a lot that came up that I had not considered previously.” Overall, Privy broadened many practitioners’ awareness of privacy risks: 10 of 24 reported either perceiving more risks or perceiving fewer benefits of their AI product concepts after using the tool.
@@
Privy encouraged participants to treat privacy as a reflective process. For instance, P9 noted that going through the AI privacy risks assessment node prompted them to refine their risk descriptions to better fit the AI product concept. Participants also emphasized that Privy made privacy assessments more enjoyable and generative, prompting deeper attention than they would normally give. As P20 explained: “it’s like the hacker mentality, right? You’re trying to break things… in a privacy context. You can just brainstorm and come up with a lot of these examples. If you’re actually doing it, you can spend like hours on this. I think it’s enjoyable to me personally.”
@@
6.3.3. Limitations of Privy-LLM and Privy-Template
While both Privy-LLM and Privy-Template effectively supported practitioners in risk envisioning and mitigation, they were less effective for the latter (Section 6.1.2). Participants’ feedback surfaced key design limitations, as well as pointed to opportunities for LLMs to more meaningfully support privacy impact assessments.

### v2

@@ 5.3. Data Analysis
We applied a mixed-methods approach to analyzing our data. Quantitative analysis. We calculated descriptive statistics (means, standard deviations) to examine how PrivyLLM and PrivyTemplate influenced practitioners’ ability to identify and mitigate privacy risks (RQ1). We fit ordinal logistic regressions to test how Privy version affected (i) privacy experts’ quality ratings of reports and (ii) practitioners’ self-reported usefulness of Privy in risk envisioning and mitigation (RQ2). Privy version was the main predictor (PrivyTemplate as reference). For expert evaluations, we included random intercepts for expert ID and, where applicable, participant ID to account for repeated measures. [^6] Specifically, for the “Per Risk” quality measure in Table 3, we included random intercepts for both expert ID and participant ID.
@@ 6. Evaluation User Study Findings
6.1. RQ1: How and to what extent do PrivyLLM and PrivyTemplate help practitioners identify and mitigate privacy risks with AI product concepts?
We first analyzed how Privy, which in both versions helps foreground privacy considerations in designing AI products, shapes practitioners’ approaches to risk identification (Section 6.1.1) and risk mitigation (Section 6.1.2). Since the findings reported in this section are shared across both PrivyLLM and PrivyTemplate, we refer to them collectively as “Privy” for simplicity.
@@ 6.1.1. Privy helps practitioners envision privacy risks that are relevant and severe
We found that Privy supported participants in identifying a broad spectrum of privacy risks for each AI product concept — practitioners surfaced at least eight different types of privacy risks, with each of those risk types being raised by two or more participants (Appendix D.4). Our findings suggest that Privy helped participants identify a diverse set of privacy risks that were both severe and relevant to their AI product concepts (DG1, DG2), from both the practitioners’ own and the privacy experts’ perspectives. Specifically, participants rated 55 (76%) of the 72 privacy risks they themselves identified as being of high relevancy and 47 (65%) of high severity. Similarly, on a 6-point Likert scale (6 as strongly agree), privacy experts, on average, rated these risks as 4.58 (SD=1.32) on relevancy and 4.97 (SD=1.07) on severity (Table 4).
@@
Privy (Overall) PrivyLLM PrivyTemplate
@@
Note. 1: strongly disagree, 2: disagree, 3: slightly disagree, 4: slightly agree, 5: agree, 6: strongly agree
@@ 6.1.2. Privy helps practitioners brainstorm risk mitigation plans that are effective and appropriate
Taken together, these results show that Privy enabled non-privacy-expert practitioners to articulate risk mitigation plans that could initiate broader privacy conversations, though additional support is needed to generate mitigation strategies that are immediately actionable and effective — a challenge we return to in Section 6.3.
This analysis also points to a notable gap in expert-assessed quality between the plans produced with PrivyLLM versus PrivyTemplate (Table 5). We explore the effects of the use of LLMs on the quality of privacy impact assessments in Section 6.2.
@@
Privy (Overall) PrivyLLM PrivyTemplate
@@
Note. 1: strongly disagree, 2: disagree, 3: slightly disagree, 4: slightly agree, 5: agree, 6: strongly agree
@@
We next examined differences between PrivyLLM and PrivyTemplate. First, we compared how privacy experts rated the quality of the privacy impact assessments produced by both PrivyLLM versus PrivyTemplate (Section 6.2.1). Then, we explored how practitioners perceived the usefulness of PrivyLLM versus PrivyTemplate in helping them complete their privacy impact assessments (Section 6.2.2). In both cases, we complemented our quantitative analysis with qualitative insights into participants’ usage patterns and opinions towards Privy’s LLM-powered features (Section 6.2.3).
6.2.1. The use of LLMs improves the quality of privacy impact assessments produced by practitioners
We evaluated the quality of privacy impact assessments produced with PrivyLLM and PrivyTemplate by analyzing the privacy risks identified and mitigation plans proposed by 24 participants (PrivyLLM: $N=12$; PrivyTemplate: $N=12$). Thirteen privacy experts independently rated these outputs on multiple quality measures (Fig. 6). Our ordinal logistic regression analyses revealed that participants using PrivyLLM produced risk statements rated significantly higher in relevancy ($\beta=1.95,p<.01$), severity ($\beta=2.13,p<.001$), correctness ($\beta=2.10,p<.001$), and clarity ($\beta=2.39,p<.001$), compared to those generated with PrivyTemplate. A similar pattern emerged for risk mitigation plans: experts judged the PrivyLLM group’s mitigation plans to be significantly more effective — “addressing identified risks” ($\beta=1.76,p<.001$) and “useful conversation starter” ($\beta=2.51,p<.001$) — and appropriate — “product specificity” ($\beta=2.04,p<.001$) and “practicality” ($\beta=1.65,p<.01$). Overall, PrivyLLM significantly outperformed PrivyTemplate across all quality measures, including overall risk identification ($\beta=2.45,p<.001$) and mitigation ($\beta=1.95,p<.001$) quality, as well as the quality of the privacy impact assessment as a whole ($\beta=2.45,p<.001$) (Fig. 7).
@@ 6.2.1. The use of LLMs improves the quality of privacy impact assessments produced by practitioners
Bar chart compares the privacy expert ratings of Privy-LLM (blue) and Privy-Template (green) on risk identification and mitigation quality, using a 6-point Likert scale (1 = strongly disagree to 6 = strongly agree, without neutral options). On the left half, it shows the quality measures of risk identification, with Privy-LLM outperforming Privy-Template in relevancy (5.09 vs. 4.06, p¡0.01), severity (5.46 vs. 4.49, p¡0.001), correctness (5.22 vs. 4.04, p¡0.001), and clarity (5.13 vs. 3.64, p¡0.001). On the right half, it shows the quality measures of proposed mitigation plans, where Privy-LLM also outperformed Privy-Template in addressing identified risks (4.37 vs. 3.35, p¡0.001), being useful conversation starters (4.69 vs. 3.42, p¡0.001), product specificity (4.96 vs. 3.46, p¡0.001), and practicality (4.42 vs. 3.42, p¡0.01).
@@
Bar chart compares the privacy expert ratings of Privy-LLM (blue) and Privy-Template (green) on overall privacy impact assessment quality, using a 6-point Likert scale (1 = strongly disagree to 6 = strongly agree, without neutral options). The Privy-LLM consistently outperformed Privy-Template across all quality measures: overall risk envisioning (4.77 vs. 2.96, p¡0.001), overall risk mitigation (4.00 vs. 2.92, p¡0.001), and overall risk impact assessment (4.42 vs. 2.88, p¡0.001).
@@ 6.2.2. The use of LLMs has minimal, if any, impact on the perceived usefulness and usability of Privy
Beyond assessing output quality, we compared PrivyLLM against PrivyTemplate on practitioners’ perceived usefulness and overall usability ratings across our five design goals (DG1-DG5). Both versions scored highly across these measures, indicating that each effectively helped users in achieving the intended goals (Fig. 8). System Usability Scale (SUS) scores were similarly strong for PrivyLLM (M=76.04, SD=15.65) and PrivyTemplate (M=75.63, SD=18.19), reflecting a “good” overall level of usability (Bangor et al., 2009).
Figure 8. Participants perceived that both PrivyLLM and PrivyTemplate were useful in achieving the design goals (DG1-DG5).
Bar chart showing participant ratings of how useful Privy-LLM (blue) and Privy-Template (green) were for achieving five design goals (DG1–DG5), measured on a 7-point Likert scale (1 = strongly disagree, 7 = strongly agree, 4 = neutral). For DG1: Help me elicit how the AI product idea entails privacy risks, Privy-LLM scored 6.17 and Privy-Template 6.50. For DG2: Help me identify privacy risks specific to this AI product idea, Privy-LLM scored 6.33 and Privy-Template 6.42. For DG3: Help me brainstorm mitigation strategies for the identified risks, Privy-LLM scored 5.67 and Privy-Template 6.17. For DG4: Create a privacy risk summary that I would use at work, Privy-LLM scored 5.00 and Privy-Template 5.50. For DG5: Feel in control when using, Privy-LLM scored 5.92 and Privy-Template 6.17.
An ordinal logistic regression showed no significant differences between versions on perceived usefulness (DG1: $\beta=-0.98,p=0.222$; DG2: $\beta=-0.15,p=0.845$; DG3: $\beta=-1.38,p=0.101$; DG4: $\beta=-0.44,p=0.566$; DG5: $\beta=-0.12,p=0.873$), and a linear regression on SUS scores likewise showed no significant difference ($\beta=0.42,p=0.953$). Taken together with Section 6.2.1, these findings suggest that while PrivyLLM supports the production of higher-quality privacy impact assessments, practitioners perceived the usability and usefulness of both versions positively — we found no significant differences between the two versions.
@@ 6.2.3. Use patterns and concerns with human-AI collaboration in privacy risk envisioning and mitigation
We next explored how participants used the LLM-powered features in PrivyLLM, and their concerns therein.
@@
Some participants used PrivyLLM as a thought partner, offering validation and reassurance. For example, as P3 put it, it served as a sanity check on their reasoning when working alone. P9 even described an “aha moment” when an LLM suggestion mirrored their own thinking: “I had come up with something without having fully read the AI suggestion that the AI thought like, this is actually something you may want to consider.”
@@
Despite the benefits of LLM support, participants raised two main concerns: overestimating the level of automated assistance provided by PrivyLLM, and the risk of overreliance on its outputs. These concerns point to a core tension: while LLMs enhanced privacy assessment quality (Section 6.2.1), they also challenged practitioners’ sense of agency in privacy work.
Some participants felt that parts of the workflow in PrivyLLM (e.g., risk mitigation) were still driven primarily by their own effort and could be improved with more LLM-assistance. As P7 noted, “I think the brainstorming mitigation tasks, I came up with them… I don’t think I took so much help from Privy.” Others described the extra cognitive effort required to integrate AI suggestions into their privacy workflow: “there’s 7–12 different things [PrivyLLM-suggested risks]… that’s a lot to juggle in my mind” (P9).
Participants also worried that LLM-generated content could bias their thinking and diminish cognitive engagement. For example, P7 made a point to generate their own ideas before consulting PrivyLLM’s provocations: “just making sure that doesn’t create bias in my head.” Similarly, P1 emphasized pausing to reflect: “have I included the risks that I felt were the most important ones before I started looking at the suggestions?” P9 further warned that less experienced practitioners may be more prone to overreliance and uncritical acceptance: “you may actually just be like, sure, this looks good, and accept.”
@@
In post-activity interviews, participants reflected on key challenges in envisioning and mitigating privacy risks for AI product concepts. We first synthesize these challenges (Section 6.3.1), drawing on the SPAF framework (Das et al., 2022) to categorize them into barriers of awareness, motivation, and ability. We then examine how Privy — across both PrivyLLM and PrivyTemplate — helped participants address some of these barriers (Section 6.3.2), and where limitations remain (Section 6.3.3).
@@
Practitioners described the absence of a systematic risk scanning mechanism, particularly when products rely on multiple data flows and third-party integrations. As P20 explained, a key challenge was “figuring out all the types of data you are taking from the user… but then you’re using some third-party solution in your product, and they’re using it, and you’re not aware of it.”
@@
Echoing prior work (Lee et al., 2024a), participants stressed that privacy is rarely treated as a core priority. P22 anticipated organizational resistance even if Privy was adopted: “I can already imagine my boss will say… this is not our first priority. We don’t need to worry about that… only if someone brings it up, or if the organization has a concern.” Some participants also described a tendency to externalize responsibility, framing privacy as the domain of specialized teams rather than part of their own remit. This distancing reduced ownership: unless privacy experts explicitly raised concerns, practitioners felt little obligation. As P6 noted, “when these external teams don’t tell us [about privacy issues], then we’re not responsible for it.”
@@
6.3.2. PrivyLLM and PrivyTemplate address major barriers
We found that both PrivyLLM and PrivyTemplate could help practitioners address these longstanding awareness, motivation, and ability barriers in privacy work.
@@ 6.3.2. PrivyLLM and PrivyTemplate address major barriers
Privy’s structured workflow provided critical scaffolding for envisioning privacy risks. Participants appreciated how the tool’s concrete definition of privacy risks and taxonomy anchored their reasoning: “understand the taxonomy here and then apply this taxonomy” (P18). Others highlighted how the workflow supported prioritization, helping them discern underlying issues. As P17 reflected, “comparing [risks] was fun… I started to realize maybe surveillance is not the big issue. Insecurity is the underlying cause.”
Privy’s structured workflow enables practitioners to ground risk envisioning from various aspects (Section 6.1.1) and to reduce blind spots (Section 6.2.3), even without LLM-powered features. As P21 noted after using PrivyTemplate: “there was a lot that came up that I had not considered previously.” Overall, Privy broadened many practitioners’ awareness of privacy risks: 10 of 24 reported either perceiving more risks or perceiving fewer benefits of their AI product concepts after using the tool.
@@
Privy encouraged participants to treat privacy as a reflective process. For instance, P9 noted that going through the AI privacy risks assessment node prompted them to refine their risk descriptions to better fit the product concept. Participants also emphasized that Privy made privacy assessments more enjoyable and generative, prompting deeper attention than they would normally give. As P20 explained: “it’s like the hacker mentality, right? You’re trying to break things… in a privacy context. You can just brainstorm and come up with a lot of these examples. If you’re actually doing it, you can spend like hours on this. I think it’s enjoyable to me personally.”
@@
6.3.3. Limitations of PrivyLLM and PrivyTemplate
While both PrivyLLM and PrivyTemplate effectively supported practitioners in risk envisioning and mitigation, they were less effective for the latter (Section 6.1.2). Participants’ feedback surfaced key design limitations, as well as pointed to opportunities for LLMs to more meaningfully support privacy impact assessments.